# An Elementary Approach to Scheduling in Generative Diffusion Models

### v1

@@ III-A KL Divergence Analysis
With the initialization $\hat{\mathbf{x}}_{t_{N}}\sim\mathcal{N}(\mathbf{0},\mathbf{I})$, time discretization sequence $\{t_{i}\}_{i=0}^{N}$, and the optimal estimator given by (5) and (7), the generated sample $\hat{\mathbf{x}}_{t_{0}}$ produced by the reverse sampling process (3) is distributed as
$p\left(\hat{\mathbf{x}}_{t_{0}}\right)=\mathcal{N}\!\left(\mathbf{0},\,\mathbf{U}\mathbf{M}\mathbf{U}^{-1}\right),\quad\mathbf{M}=\mathrm{diag}(m_{1},\ldots,m_{k}),$ (8)
@@
By (6), we can rewrite $q(\mathbf{x}_{t_{0}})=\mathcal{N}(\mathbf{0},\mathbf{U}\mathbf{N}\mathbf{U}^{-1})$, where $\mathbf{N}=\mathrm{diag}(n_{1},\ldots,n_{k})$ with
@@
$D_{\mathrm{KL}}\!\left(p(\hat{\mathbf{x}}_{t_{0}})\,\|\,q(\mathbf{x}_{t_{0}})\right)=\frac{1}{2}\sum_{\ell=1}^{k}\left(\frac{m_{\ell}}{n_{\ell}}-\log\frac{m_{\ell}}{n_{\ell}}-1\right),$ (11)
@@
$D_{\mathrm{KL}}\!\left(p(\hat{\mathbf{x}}_{t_{0}})\,\|\,q(\mathbf{x}_{t_{0}})\right)$ $=\frac{1}{2}\sum_{\ell=1}^{k}\bigl(e^{2r_{\ell}}-2r_{\ell}-1\bigr)$
@@
From (18), we observe that the KL divergence scales as $\mathcal{O}(1/N^{2})$. As $N\to\infty$, the eigenvalue mismatch vanishes (i.e., $m_{\ell}/n_{\ell}\to 1$), leading to $D_{\mathrm{KL}}(p(\hat{\mathbf{x}}_{t_{0}})\|q(\mathbf{x}_{t_{0}}))\to 0$. This verifies that the reverse sampling process in (3) is asymptotically consistent, recovering the exact source distribution in the continuous-time limit, in accordance with the classical theory well known for DMs (see, e.g., [1]).
III-B Derivation of Tangent Law
@@
$\text{subject to }\eta(0)=0\text{ and }\lim_{t\to 1}\eta(t)=\infty.$ (21)
@@
Given the boundary conditions $\eta(0)=0$ and $\lim_{t\to 1}\eta(t)=\infty$, the transformed boundary values are $Q_{\ell}(0)=0$ and $Q_{\ell}(1)=\frac{\pi}{2\sqrt{\mu_{\ell}}}$. Therefore, the optimal trajectory is $Q_{\ell}(t)=\frac{\pi t}{2\sqrt{\mu_{\ell}}}$. Finally, combining the definition of $Q_{\ell}(t)$ (22) yields the tangent law schedule $\eta_{\ell}(t)=\sqrt{\mu_{\ell}}\tan(\frac{\pi}{2}t)$. ∎
@@
III-C Time Discretization Strategy Evaluation
@@
To apply (11), we model the source distribution using a synthetic power-law covariance spectrum, which mimics the heavy-tailed eigenvalue decay characteristic of real image data [21]. Figure 1 presents the quantitative evaluation of the closed-form KL divergence (11) across varying step budgets $N$ (details in Appendix E). This comparison covers both VP and VE settings as considered in [20] with the same time discretization strategy given in (30). The results demonstrate that in both settings, the discretization strategy (30) with an appropriate choice of $\rho$ can achieve notably faster convergence compared to the widely used uniform-$\lambda$ baseline, significantly reducing the reverse sampling mismatch under limited budgets. These preliminary observations motivate further experimentation on real-world image generation tasks, as presented in the next section.
IV Experiments
@@ Appendix A Proof of Lemma 1
$\hat{\mathbf{x}}_{t_{N}}\sim\mathcal{N}(\mathbf{0},\sigma_{t_{N}}^{2}\mathbf{I}),\quad\text{where}\quad\sigma_{t_{N}}=\begin{cases}1&\text{VP setting},\\ \sigma_{\max}&\text{VE setting}.\end{cases}$ (37)
@@
$p(\hat{\mathbf{x}}_{t_{0}})=\mathcal{N}\!\left(\mathbf{0},\,\mathbf{U}\mathbf{M}\mathbf{U}^{-1}\right),$ (38)
@@
$\mathbf{M}$ $\triangleq\sigma_{t_{N}}^{2}\left(\prod_{j=1}^{N}\mathbf{D}_{j}\right)\left(\prod_{j=1}^{N}\mathbf{D}_{j}\right)^{\!\top}=\mathrm{diag}(m_{1},\ldots,m_{k}),$ (39)
@@
$m_{\ell}=\sigma_{t_{N}}^{2}\prod_{j=1}^{N}d_{j,\ell}^{2},\qquad\ell=1,\ldots,k,$ (40)
@@
$D_{\mathrm{KL}}\!\left(p(\hat{\mathbf{x}}_{t_{0}})\,\|\,q(\mathbf{x}_{t_{0}})\right)$ $=D_{\mathrm{KL}}\!\left(\mathcal{N}\!\left(\mathbf{0},\,\mathbf{U}\mathbf{M}\mathbf{U}^{-1}\right)\,\|\,\mathcal{N}\!\left(\mathbf{0},\,\mathbf{U}\mathbf{N}\mathbf{U}^{-1}\right)\right)$
@@ B-A Closed Form of Integral Term $I_{\ell}$
To interpret this result, we first recall that $n_{\ell}\triangleq\alpha^{2}(0)\mu_{\ell}+\sigma^{2}(0)$ (10). We now consider the following two boundary conditions, as we mentioned before in Section II-A.
  • Variance Preserving (VP): The boundary conditions are specified as $\alpha(0)=1,\sigma(0)=0$ and $\alpha(1)=0,\sigma(1)=1$. Substituting these into (61) yields
@@
  • Variance Exploding (VE): The boundary conditions are $\alpha(t)\equiv 1$, $\sigma(0)=0$, and $\sigma(1)=\sigma_{\max}$. In this case, the integral (61) becomes
@@ Appendix C Supplementary Details of Theorem 1
Given the noise schedule $\alpha(t)\equiv 1$, $\sigma(0)=0$, and $\sigma(1)=\sigma_{\max}$ for the VE setting, along with the definition $\eta(t)=\sigma(t)/\alpha(t)$, we derive the boundary conditions as follows:
$\eta(0)=0\quad\text{and}\quad\eta(1)=\sigma_{\max}.$ (71)
@@ D-A Derivation of Leading-Order Coefficient
  • Variance Preserving (VP): Since $\alpha(1)=0$ and $\sigma(1)=1$, we have $\eta_{1}\to\infty$. Consequently, the angle becomes
@@
  • Variance Exploding (VE): Since $\alpha(t)\equiv 1$ and $\sigma(1)=\sigma_{\max}$, we have $\eta_{1}=\sigma_{\max}$,
@@
Same as we mentioned before in Appendix B-A and Appendix C, under the condition of $\sigma_{\max}^{2}\gg\gamma$, the VP setting solution can be regarded as an accurate approximation of the VE setting.
@@
The standard integral result from [2, formula 2.563.1 and 2.562.1 in pp.177] is
@@ D-B Numerical Results on Theorem 2
To assess the performance of the $\gamma^{\star}$ parameterized tangent law noise schedule derived in Theorem 2, we conduct a two-fold evaluation based on the closed-form KL divergence (11), with results summarized in Figure 2. First, we compare our parameterized tangent law schedule (using $\gamma^{\star}$) against representative noise schedules widely employed in DMs (Figure 2 top row). The results indicate that our method achieves faster convergence in both low-dimensional and high-dimensional synthetic Gaussian sources (details in Appendix E). Second, we perform an ablation study on the hyperparameter $\gamma$ (Figure 2 bottom row). The empirical results align with our theoretical analysis, showing that the derived $\gamma^{\star}$ (26) leads to the best performance compared to other choices.
@@
Appendix E Implementation Details

### v2

@@ IV-A KL Divergence Analysis
With the initialization $\hat{\mathbf{x}}_{t_{N}}\sim\mathcal{N}(\mathbf{0},\mathbf{I})$, time discretization sequence $\{t_{i}\}_{i=0}^{N}$, and the optimal estimator given by (7) and (22), the generated sample $\hat{\mathbf{x}}_{t_{0}}$ produced by deterministic reverse sampling (10) is distributed as
$\hat{p}_{0,G}\left(\hat{\mathbf{x}}_{t_{0}}\right)=\mathcal{N}\!\left(\mathbf{0},\,\mathbf{U}\mathbf{M}\mathbf{U}^{-1}\right),\quad\mathbf{M}=\mathrm{diag}(m_{1},\ldots,m_{k}),$ (23)
@@
By (21), we can rewrite $q_{0,G}(\mathbf{x}_{t_{0}})=\mathcal{N}(\mathbf{0},\mathbf{U}\mathbf{N}\mathbf{U}^{-1})$, where $\mathbf{N}=\mathrm{diag}(n_{1},\ldots,n_{k})$ with
@@
$D_{\mathrm{KL}}\!\left(\hat{p}_{0,G}(\hat{\mathbf{x}}_{t_{0}})\,\|\,q_{0,G}(\mathbf{x}_{t_{0}})\right)=\frac{1}{2}\sum_{\ell=1}^{k}\left(\frac{m_{\ell}}{n_{\ell}}-\log\frac{m_{\ell}}{n_{\ell}}-1\right),$ (26)
@@
$D_{\mathrm{KL}}\!\left(\hat{p}_{0,G}(\hat{\mathbf{x}}_{t_{0}})\,\|\,q_{0,G}(\mathbf{x}_{t_{0}})\right)$ $=\frac{1}{2}\sum_{\ell=1}^{k}\bigl(e^{2r_{\ell}}-2r_{\ell}-1\bigr)$
@@
From (33), we observe that the KL divergence scales as $\mathcal{O}(1/N^{2})$. As $N\to\infty$, the eigenvalue mismatch vanishes (i.e., $m_{\ell}/n_{\ell}\to 1$), leading to $D_{\mathrm{KL}}(p_{0,G}(\hat{\mathbf{x}}_{t_{0}})\|q_{0,G}(\mathbf{x}_{t_{0}}))\to 0$. This verifies that the reverse sampling process in (10) is asymptotically consistent, recovering the exact source distribution in the continuous-time limit, in accordance with the classical theory well known for DMs (see, e.g., [1]).
IV-B Derivation of Tangent Law
@@
$\text{subject to }\lim_{t\to 0}\eta(t)=0\text{ and }\lim_{t\to 1}\eta(t)=\infty.$ (36)
@@
Given the boundary conditions in (36), the transformed boundary values are $Q_{\ell}(0)=0$ and $Q_{\ell}(1)=\frac{\pi}{2\sqrt{\mu_{\ell}}}$. Therefore, the optimal trajectory is $Q_{\ell}(t)=\frac{\pi t}{2\sqrt{\mu_{\ell}}}$. Finally, combining the definition of $Q_{\ell}(t)$ (37) yields the tangent law schedule $\eta_{\ell}(t)=\sqrt{\mu_{\ell}}\tan(\frac{\pi}{2}t)$. ∎
@@
IV-C Time Discretization Strategy Evaluation
@@
To apply (26), we model the source distribution using a synthetic power-law covariance spectrum, which mimics the heavy-tailed eigenvalue decay characteristic of real image data [26]. Figure 1 presents the quantitative evaluation of the closed-form KL divergence (26) across varying step budgets $N$ (details in Appendix F). This comparison covers both VP and VE settings as considered in [25] with the same time discretization strategy given in (45). The results demonstrate that in both settings, the discretization strategy (45) with an appropriate choice of $\rho$ can achieve notably faster convergence compared to the widely used uniform-$\lambda$ baseline, significantly reducing the reverse sampling mismatch under limited budgets. These preliminary observations motivate further experimentation on real-world image generation tasks, as presented in the next section.
V Experiments
@@ Appendix A Proof of Lemma 1
$\hat{\mathbf{x}}_{t_{N}}\sim\mathcal{N}(\mathbf{0},\sigma_{N}^{2}\mathbf{I}),\quad\text{where}\quad\sigma_{N}=\begin{cases}1&\text{VP setting},\\ \sigma_{\max}&\text{VE setting}.\end{cases}$ (52)
@@
$\hat{p}_{0,G}(\hat{\mathbf{x}}_{t_{0}})=\mathcal{N}\!\left(\mathbf{0},\,\mathbf{U}\mathbf{M}\mathbf{U}^{-1}\right),$ (53)
@@
$\mathbf{M}$ $\triangleq\sigma_{N}^{2}\left(\prod_{j=1}^{N}\mathbf{D}_{j}\right)\left(\prod_{j=1}^{N}\mathbf{D}_{j}\right)^{\!\top}=\mathrm{diag}(m_{1},\ldots,m_{k}),$ (54)
@@
$m_{\ell}=\sigma_{N}^{2}\prod_{j=1}^{N}d_{j,\ell}^{2},\qquad\ell=1,\ldots,k,$ (55)
@@
$D_{\mathrm{KL}}\!\left(\hat{p}_{0,G}(\hat{\mathbf{x}}_{t_{0}})\,\|\,q_{0,G}(\mathbf{x}_{t_{0}})\right)$ $=D_{\mathrm{KL}}\!\left(\mathcal{N}\!\left(\mathbf{0},\,\mathbf{U}\mathbf{M}\mathbf{U}^{-1}\right)\,\|\,\mathcal{N}\!\left(\mathbf{0},\,\mathbf{U}\mathbf{N}\mathbf{U}^{-1}\right)\right)$
@@ B-A Closed Form of Integral Term $I_{\ell}$
To interpret this result, we first recall that $n_{\ell}\triangleq\alpha^{2}(0)\mu_{\ell}+\sigma^{2}(0)$ (25). We now consider the following two boundary conditions, as we mentioned before in Section II.
  • Variance Preserving (VP): The boundary conditions are specified as $\lim_{t\to 0}\alpha(t)=1$, $\lim_{t\to 0}\sigma(t)=0$, and typically $\lim_{t\to 1}\alpha(t)=0$, $\lim_{t\to 1}\sigma(t)=1$. Substituting these into (76) yields
@@
  • Variance Exploding (VE): The boundary conditions are $\alpha(t)\equiv 1$, $\lim_{t\to 0}\sigma(t)=0$, and $\lim_{t\to 1}\sigma(t)=\sigma_{\max}$. In this case, the integral (76) becomes
@@ Appendix C Supplementary Details of Theorem 2
Given the noise schedule $\alpha(t)\equiv 1$, $\lim_{t\to 0}\sigma(t)=0$, and $\lim_{t\to 1}\sigma(t)=\sigma_{\max}$ for the VE setting, along with the definition $\eta(t)=\sigma(t)/\alpha(t)$, we derive the boundary conditions as follows:
$\lim_{t\to 0}\eta(t)=0\quad\text{and}\quad\lim_{t\to 1}\eta(t)=\sigma_{\max}.$ (86)
@@ D-A Derivation of Leading-Order Coefficient
  • Variance Preserving (VP): Since $\lim_{t\to 1}\alpha(t)=0$, $\lim_{t\to 1}\sigma(t)=1$, we have $\eta_{1}\to\infty$. Consequently, the angle becomes
@@
  • Variance Exploding (VE): Since $\alpha(t)\equiv 1$ and $\lim_{t\to 1}\sigma(t)=\sigma_{\max}$, we have $\eta_{1}=\sigma_{\max}$,
@@
Same as we stated before in Appendix B-A and Appendix C, under the condition of $\sigma_{\max}^{2}\gg\gamma$, the VP setting solution can be regarded as an accurate approximation of the VE setting.
@@
The standard integral result from [5, formula 2.563.1 and 2.562.1 in pp.177] is
@@ D-B Numerical Results on Theorem 3
To assess the performance of the $\gamma^{\star}$ parameterized tangent law noise schedule derived in Theorem 3, we conduct a two-fold evaluation based on the closed-form KL divergence (26), with results summarized in Figure 2. First, we compare our parameterized tangent law schedule (using $\gamma^{\star}$) against representative noise schedules widely employed in DMs (Figure 2 top row). The results indicate that our method achieves faster convergence in both low-dimensional and high-dimensional synthetic Gaussian sources (details in Appendix F). Second, we perform an ablation study on the hyperparameter $\gamma$ (Figure 2 bottom row). The empirical results align with our theoretical analysis, showing that the derived $\gamma^{\star}$ (41) leads to the best performance compared to other choices.
@@
Appendix E Lemma of Girsanov’s Theorem
Lemma 3.
Let $\mathbb{P}_{\theta}$ denote the path measure induced by the continuous-time SDE parameterized by the neural network $\bm{\epsilon}_{\theta}$, and let $\tilde{\mathbb{P}}_{\theta}$ denote the measure induced by its discretization scheme:
$\mathbb{P}_{\theta}:\quad$ $\mathrm{d}\mathbf{x}_{t}=\left(f_{t}\mathbf{x}_{t}-h_{t}\bm{\epsilon}_{\theta}(\mathbf{x}_{t},t)\right)\mathrm{d}t+g_{t}\mathrm{d}\bar{\mathbf{w}}_{t},\quad\mathbf{x}_{1}\sim q_{1}(\mathbf{x}_{1}),$ (96a)
$\tilde{\mathbb{P}}_{\theta}:\quad$ $\mathrm{d}\mathbf{x}_{t}=\left(f_{t}\mathbf{x}_{t}-h_{t}\bm{\epsilon}_{\theta}(\mathbf{x}_{t_{j}},t_{j})\right)\mathrm{d}t+g_{t}\mathrm{d}\bar{\mathbf{w}}_{t},\quad\mathbf{x}_{1}\sim\mathcal{N}(\mathbf{0},\sigma_{N}^{2}\mathbf{I}).$ (96b)
The KL divergence between these two processes satisfies:
$D_{KL}(\mathbb{P}_{\theta}||\tilde{\mathbb{P}}_{\theta})$ $=D_{KL}(q_{1}(\mathbf{x}_{1})||\mathcal{N}(\mathbf{0},\sigma_{N}^{2}\mathbf{I}))+\frac{1}{2}\int_{0}^{1}\frac{h_{t}^{2}}{g_{t}^{2}}\,\mathbb{E}_{\mathbf{x}_{t}\sim p_{t}}\left[\|\bm{\epsilon}_{\theta}(\mathbf{x}_{t},t)-\bm{\epsilon}_{\theta}(\mathbf{x}_{t_{j}},t_{j})\|^{2}\right]\mathrm{d}t$ (97)
$=D_{KL}(q_{1}(\mathbf{x}_{1})||\mathcal{N}(\mathbf{0},\sigma_{N}^{2}\mathbf{I}))+\sum_{j=1}^{N}\frac{1}{4}h_{t_{j}}^{2}\delta_{j}^{2}\,\mathbb{E}_{\mathbf{x}_{t_{j}}\sim p_{t_{j}}}\left[\|\mathbf{J}_{\bm{\epsilon}}(\mathbf{x}_{t_{j}},t_{j})\|_{F}^{2}\right]+\mathcal{R}(\delta),$ (98)
where $\delta_{j}=t_{j}-t_{j-1}$ is the step size, $\mathbf{J}_{\bm{\epsilon}}\triangleq\nabla_{\mathbf{x}}\bm{\epsilon}_{\theta}(\mathbf{x},t)$ denotes the Jacobian matrix of $\bm{\epsilon}_{\theta}$, and $\mathcal{R}(\delta)$ represents the cumulative higher-order error terms satisfying $\mathcal{R}(\delta)=\mathcal{O}(\delta^{2})$ as $\delta\triangleq\max_{j}\delta_{j}\to 0$.
Proof.
The proof proceeds in two steps. First, we apply Girsanov’s theorem to derive the exact integral form of the KL divergence. Second, we analyze the local discretization error using Itô’s calculus to obtain the leading-order expansion.
To facilitate the rigorous derivation of the error bound, we first rely on the following regularity conditions. These assumptions are standard in the theoretical analysis of DMs:
  • Condition 1 (Coefficients): The SDE coefficients $f_{t}$ and $g_{t}$ are Lipschitz continuous with respect to time $t$ and bounded on $[0,1]$.
  • Condition 2 (Neural Network): The neural estimator $\bm{\epsilon}_{\theta}(\mathbf{x},t)$ is twice continuously differentiable ($C^{2}$) with respect to $\mathbf{x}$ and $C^{1}$ with respect to $t$. Moreover, both the network output and its Jacobian are bounded.
  • Condition 3 (Bounded Moments): The drift term of the score evolution (defined later in Eq. (E)) possesses a bounded second moment, i.e., $\sup_{t\in[0,1]}\mathbb{E}[\|\bm{\phi}(\mathbf{x}_{t},t)\|^{2}]<\infty$.
Let $u_{t}$ denote the difference between the drift terms of $\mathbb{P}_{\theta}$ and $\hat{\mathbb{P}}_{\theta}$, normalized by the diffusion coefficient $g_{t}$. We define:
$u_{t}\triangleq\frac{-h_{t}\left(\bm{\epsilon}_{\theta}(\mathbf{x}_{t},t)-\bm{\epsilon}_{\theta}(\mathbf{x}_{t_{j}},t_{j})\right)}{g_{t}}.$ (99)
To ensure the validity of the measure change, we assume Novikov’s condition holds, which requires that the following expectation is finite:
$\mathbb{E}\left[\exp\left(\frac{1}{2}\int_{0}^{1}\|u_{t}\|^{2}\mathrm{d}t\right)\right]=\mathbb{E}\left[\exp\left(\frac{1}{2}\int_{0}^{1}\frac{h_{t}^{2}}{g_{t}^{2}}\left\|\bm{\epsilon}_{\theta}(\mathbf{x}_{t},t)-\bm{\epsilon}_{\theta}(\mathbf{x}_{t_{j}},t_{j})\right\|^{2}\mathrm{d}t\right)\right]<\infty.$ (100)
Under conditions 1 and 2, Girsanov’s theorem [19] applies. The KL divergence decomposes into the divergence of the initial distributions and the expected path integral of the squared norm of $u_{t}$:
$D_{KL}(\mathbb{P}_{\theta}||\tilde{\mathbb{P}}_{\theta})=D_{KL}(q_{1}||\mathcal{N}(\mathbf{0},\sigma_{N}^{2}\mathbf{I}))+\frac{1}{2}\int_{0}^{1}\mathbb{E}\left[\|u_{t}\|^{2}\right]\mathrm{d}t.$ (101)
Substituting the definition of $u_{t}$ back into Eq. (101) yields the equality given in (97).
We focus on the expectation term within the interval $t\in(t_{j-1},t_{j}]$. Applying Itô’s Lemma [19] to $\bm{\epsilon}_{\theta}(\mathbf{x}_{t},t)$, its dynamics follow:
$\mathrm{d}\bm{\epsilon}_{\theta}(\mathbf{x}_{t},t)$ $=\left(\frac{\partial\bm{\epsilon}_{\theta}}{\partial t}+\mathbf{J}_{\bm{\epsilon}}(\mathbf{x}_{t},t)\left(f_{t}\mathbf{x}_{t}-h_{t}\bm{\epsilon}_{\theta}(\mathbf{x}_{t},t)\right)+\frac{1}{2}g_{t}^{2}\nabla_{\mathbf{x}}^{2}\bm{\epsilon}_{\theta}(\mathbf{x}_{t},t)\right)\mathrm{d}t+g_{t}\mathbf{J}_{\bm{\epsilon}}(\mathbf{x}_{t},t)\mathrm{d}\bar{\mathbf{w}}_{t}$
$\triangleq\bm{\phi}(\mathbf{x}_{t},t)\mathrm{d}t+g_{t}\mathbf{J}_{\bm{\epsilon}}(\mathbf{x}_{t},t)\mathrm{d}\bar{\mathbf{w}}_{t},$ (102)
where $\mathbf{J}_{\bm{\epsilon}}=\nabla_{\mathbf{x}}\bm{\epsilon}_{\theta}\in\mathbb{R}^{d\times d}$ is the Jacobian matrix and $\nabla_{\mathbf{x}}^{2}$ denotes the vector Laplacian operator. Integrating from $t$ to $t_{j}$:
$\bm{\epsilon}_{\theta}(\mathbf{x}_{t_{j}},t_{j})-\bm{\epsilon}_{\theta}(\mathbf{x}_{t},t)=\int_{t}^{t_{j}}\bm{\phi}(\mathbf{x}_{\tau},\tau)\mathrm{d}\tau+\int_{t}^{t_{j}}g_{\tau}\mathbf{J}_{\bm{\epsilon}}(\mathbf{x}_{\tau},\tau)\mathrm{d}\bar{\mathbf{w}}_{\tau}.$ (103)
Then utilizing the Itô Isometry property [19], we get
$\mathbb{E}\left[\|\bm{\epsilon}_{\theta}(\mathbf{x}_{t},t)-\bm{\epsilon}_{\theta}(\mathbf{x}_{t_{j}},t_{j})\|^{2}\right]$ $=\int_{t}^{t_{j}}g_{\tau}^{2}\mathbb{E}\left[\|\mathbf{J}_{\bm{\epsilon}}(\mathbf{x}_{\tau},\tau)\|_{F}^{2}\right]\mathrm{d}\tau+\mathbb{E}\left[\left\|\int_{t}^{t_{j}}\bm{\phi}(\mathbf{x}_{\tau},\tau)\mathrm{d}\tau\right\|^{2}\right].$ (104)
Assuming the time discretization is sufficiently fine (i.e., the step size $\delta_{j}\triangleq t_{j}-t_{j-1}$ is small enough), we can apply local expansions based on Conditions 1, 2, and 3:
  • Using the Lipschitz continuity of the coefficients (Condition 1), we approximate the ratio $\frac{h_{t}^{2}}{g_{t}^{2}}$ by its value at $t_{j}$: $\frac{h_{t}^{2}}{g_{t}^{2}}=\frac{h_{t_{j}}^{2}}{g_{t_{j}}^{2}}+\mathcal{O}(t_{j}-t).$
  • Let $\Phi(\tau)\triangleq g_{\tau}^{2}\mathbb{E}[\|\mathbf{J}_{\bm{\epsilon}}(\mathbf{x}_{\tau},\tau)\|_{F}^{2}]$. Under Conditions 1 and 2, $\Phi(\tau)$ is Lipschitz continuous. Expanding around $t_{j}$: $\Phi(\tau)=\Phi(t_{j})+\mathcal{O}(t_{j}-\tau)$.
  • Since the drift $\bm{\phi}$ has a bounded second moment (Condition 3), applying the Cauchy-Schwarz inequality yields:
    $\mathbb{E}\left[\left\|\int_{t}^{t_{j}}\bm{\phi}(\mathbf{x}_{\tau},\tau)\mathrm{d}\tau\right\|^{2}\right]\leq(t_{j}-t)\int_{t}^{t_{j}}\mathbb{E}\left[\|\bm{\phi}\|^{2}\right]\mathrm{d}\tau=\mathcal{O}((t_{j}-t)^{2}).$ (105)
Substituting these estimates back into the local error expectation (104), we obtain
$\mathbb{E}\left[\|\bm{\epsilon}_{\theta}(\mathbf{x}_{t},t)-\bm{\epsilon}_{\theta}(\mathbf{x}_{t_{j}},t_{j})\|^{2}\right]$ $=\int_{t}^{t_{j}}\left(\Phi(t_{j})+\mathcal{O}(t_{j}-\tau)\right)\mathrm{d}\tau+\mathcal{O}((t_{j}-t)^{2})$
$=\Phi(t_{j})(t_{j}-t)+\int_{t}^{t_{j}}\mathcal{O}(t_{j}-\tau)\mathrm{d}\tau+\mathcal{O}((t_{j}-t)^{2})$
$=g_{t_{j}}^{2}\mathbb{E}_{\mathbf{x}_{t_{j}}}[\|\mathbf{J}_{\bm{\epsilon}}(\mathbf{x}_{t_{j}},t_{j})\|_{F}^{2}](t_{j}-t)+\mathcal{O}((t_{j}-t)^{2}).$ (106)
Now, we substitute Eq. (E) back into the integral term given in (101). Focusing on the contribution from the $j$-th interval $(t_{j-1},t_{j}]$, the term involving $u_{t}$ becomes
$\int_{t_{j-1}}^{t_{j}}\mathbb{E}\left[\|u_{t}\|^{2}\right]\mathrm{d}t$ $=\int_{t_{j-1}}^{t_{j}}\frac{h_{t}^{2}}{g_{t}^{2}}\mathbb{E}\left[\|\bm{\epsilon}_{\theta}(\mathbf{x}_{t},t)-\bm{\epsilon}_{\theta}(\mathbf{x}_{t_{j}},t_{j})\|^{2}\right]\mathrm{d}t$
$=\int_{t_{j-1}}^{t_{j}}\left(\frac{h_{t_{j}}^{2}}{g_{t_{j}}^{2}}+\mathcal{O}(t_{j}-t)\right)\left(g_{t_{j}}^{2}\mathbb{E}[\|\mathbf{J}_{\bm{\epsilon}}(\mathbf{x}_{t_{j}},t_{j})\|_{F}^{2}](t_{j}-t)+\mathcal{O}((t_{j}-t)^{2})\right)\mathrm{d}t$
$=\int_{t_{j-1}}^{t_{j}}\left(h_{t_{j}}^{2}\mathbb{E}[\|\mathbf{J}_{\bm{\epsilon}}(\mathbf{x}_{t_{j}},t_{j})\|_{F}^{2}](t_{j}-t)+\mathcal{O}((t_{j}-t)^{2})\right)\mathrm{d}t.$ (107)
So we have the integral of the squared drift difference
$\int_{0}^{1}\mathbb{E}\left[\|u_{t}\|^{2}\right]\mathrm{d}t$ $=\sum_{j=1}^{N}\int_{t_{j-1}}^{t_{j}}\mathbb{E}\left[\|u_{t}\|^{2}\right]\mathrm{d}t$
$=\sum_{j=1}^{N}\left(h_{t_{j}}^{2}\mathbb{E}[\|\mathbf{J}_{\bm{\epsilon}}(\mathbf{x}_{t_{j}},t_{j})\|_{F}^{2}]\int_{t_{j-1}}^{t_{j}}(t_{j}-t)\mathrm{d}t+\int_{t_{j-1}}^{t_{j}}\mathcal{O}((t_{j}-t)^{2})\mathrm{d}t\right)$
$=\sum_{j=1}^{N}\left(\frac{1}{2}h_{t_{j}}^{2}\delta_{j}^{2}\mathbb{E}[\|\mathbf{J}_{\bm{\epsilon}}(\mathbf{x}_{t_{j}},t_{j})\|_{F}^{2}]+\mathcal{O}(\delta_{j}^{3})\right).$ (108)
Finally, substituting this back into the Girsanov formula (101), and summing over all intervals $j=1,\dots,N$
$D_{KL}(\mathbb{P}_{\theta}||\tilde{\mathbb{P}}_{\theta})$ $=D_{KL}(q_{1}||\mathcal{N}(\mathbf{0},\sigma_{N}^{2}\mathbf{I}))+\frac{1}{2}\sum_{j=1}^{N}\left(\frac{1}{2}h_{t_{j}}^{2}\delta_{j}^{2}\mathbb{E}[\|\mathbf{J}_{\bm{\epsilon}}\|_{F}^{2}]+\mathcal{O}(\delta_{j}^{3})\right)$
$=D_{KL}(q_{1}||\mathcal{N}(\mathbf{0},\sigma_{N}^{2}\mathbf{I}))+\sum_{j=1}^{N}\frac{1}{4}h_{t_{j}}^{2}\delta_{j}^{2}\mathbb{E}[\|\mathbf{J}_{\bm{\epsilon}}\|_{F}^{2}]+\mathcal{R}(\delta),$ (109)
where $\mathcal{R}(\delta)$ represents the higher-order terms, which satisfies $\mathcal{R}(\delta)\triangleq\mathcal{O}((\max_{j}\delta_{j})^{2})$ as $\delta=\max_{j}\delta_{j}\to 0$. This recovers the exact form stated in (98). ∎
Appendix F Implementation Details